TROYBOY1:

how do I find x+y, if: 5x+8y=67 and 2x-y=31. a. 98. b. 46. c. 21. d. 14. e. 7. my teacher asked me too find this but im a little confused where do i distribute and evaluate

fwval:

first we do equation 2x - 7 = 31 rearrange to isolate y = 2x - 31 substitute y into 1 5x + 8 (2x - 31) = 67 simplify 5x + 16x - 248 = 67 combine like terms 21x - 248 = 67 solve for x add 248 both sides 21x = 315 divide by 21 x = 15 then we find y y = 2x - 31 substitute x = 15 y = 2 (15) - 31 y = 30 - 31 y = -1 then we calculate x + y add x and y x + y = 15 + (-1) = 14 answer for x + y = 14 so the answer is d

fwval:

is that what u needed?

TROYBOY1:

@fwval wrote:
is that what u needed?
YES! TY

fwval:

@troyboy1 wrote:
@fwval wrote:
is that what u needed?
YES! TY
yw

surjithayer:

5x+8y=67 2x-y=31 adding both 7x+7y=98 divide by 7 x+y=14